Helen is now 40 years old and arlene is 20 years old.How many years ago was Helen three times as old as Arlene was then.
Answer by Paul(988)

Helen is 40 years old
Arlene is 20 years old.
TO determine: thrice the age of Arlene.
Eqaution:
3(20-x)=40-x
-3x+60=40-x
-2x=-20
x=10
Hence, 10 years ago, she was thrice as old as her.
Paul.